A resort owner wants to enclose a beachfront area for swimming activities. Based on her plan, only 3 sides will be fence with 270 meter rope and floats, while the shoreline part will be open. Determine the dimension of the 3 sides of the rectangle that will give a maximum area.

1
Expert's answer
2021-11-15T16:44:08-0500

Lets draw a picture



We have to find such value 0 < x < 270 that maximize the area of the blue rectangle. Let S be the area, then "S(x)=x*(270-2x)=270x-2x^2"

Lets find the critical points

"S'(x)=270-4x"

"270-4x=0\implies x =67.5". Since to the left from x = 67.5 the value of S'(X) is positive, and to the right it's negative, then x = 67.5 is the point of maximum

The sides of a rectangle is 67.5 and 270 - 2 * 67.5 = 135 meters and the area is"S(67.5)=270*67.5-2*67.5^2=9112.5" square meters